solve by the substitution method. 5x+9y=51 and x=39-5y
Answer by richwmiller(17219)

5*(39-5y)+9y=51
multiply out and solve for y
5*39-25y+9y=51
195-16y=51
144=16y
4*4*3*3=4*4y
9=y
x=39-5y
x=39-45
x=-6